OpenStudy (anonymous):

Perform the indicated operation. 13/16 - 3/8 I got 56/128 in all, but it needs to be lowered as low as it can get. How would I do this?

OpenStudy (campbell_st):

3/8 = 6/16 so the problem is 13/16 - 6/16 after using an equivalent fraction and common denominators which can be written \[\frac{13 - 6}{16}\] just complete the subtraction in the numerator for the answer

OpenStudy (campbell_st):

as a fraction you need 13 - 6 = 7 so the answer is 7/16

OpenStudy (anonymous):

How did you get the 6?

OpenStudy (campbell_st):

3/8 double the top and bottom of the fraction gives 6/16 this is called an equivalent fraction. That ,means it has the same value as 3/8 but can be written in a different way